OpenStudy (anonymous):

PreCalc help?! Given f(3)=80 and f(6)=120, find f(24) for f(x) as: a. linear function b. power function c. exponential function

OpenStudy (haseeb96):

Exponential function

OpenStudy (anonymous):

a) f(3)=80 -> (x,y)=(3,80) f(6)=120 -> (x,y) = (6,120) find slope: (120-80)/(6-3)= 40/3 use y=mx+b to figure out b: 80=(40/3)(3)+b 80=40+b 40=b .... so f(x)= (40/3)x+40 and f(24)= (40/30)(24)= 360

OpenStudy (anonymous):

Thank you for the explaination; I have no idea how to do this stuff. Any idea on how to do the others?

OpenStudy (anonymous):

b)\[ f(x)=ax^b\\ 80=a(3)^b\\ 120=a(6)^b \]

OpenStudy (anonymous):

So what do I do after that? Do I divide by 6 on both sides?

OpenStudy (anonymous):

Divide first equation by second equation: \[ \frac 23=\left(\frac 36\right)^b \]Notice that \(a\)s cancel.

OpenStudy (anonymous):

\[ \ln 2/3=b\ln 3/6 \]

OpenStudy (anonymous):

Do I put 24 in for b?

OpenStudy (anonymous):

No, \[ b=\frac{\ln 2/3}{\ln 1/2} \]

OpenStudy (anonymous):

You can use a calculator to approximate I guess.

OpenStudy (anonymous):

Alright, I'll do that. Thanks!

OpenStudy (anonymous):

Then use \[ 80=a(3)^b \]To solve for \(a\)